# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=gray!40

### v1

@@ A.3.2 Deploying DistR
Following configuration, build DistR as follows: {mdframed}[style=MyFrame,nobreak=true] # Compile the communication static library
@@
# Copy the static library to the DistR directory
@@
cargo build --release
@@ A.3.3 Running Applications
DistR is bundled with four example applications: Dataframe, GEMM, KVStore, and SocialNet. Follow these steps to execute them:
@@
    Launch the DistR executable on all servers, excluding the main server: {mdframed}[style=MyFrame,nobreak=true] # Start the DistR process with the specified server index and application name.
@@
    On the main server: {mdframed}[style=MyFrame,nobreak=true] # Start the main DistR process with the specified application.
@@
More details of DistR’s installation and deployment can be found in DistR’s code repository.
@@ B.1 Pointer Layout
DistR extends each Rust Box pointer and reference with an additional 64-bit field, as depicted in Figure 8. Additionally, in the global address field, we reserve 16 bits as “color” bits, which are used to track the version of the referenced value. Every time a mutable reference is dereferenced for the first time, or the owner performs a write operation subsequent to the return of all immutable references, the color bits are updated. Consequently, this mechanism guarantees that any newly created immutable references will obtain a new global address [^3] In the following text, a “colored global address” means the composition of a global address and its color.. This update is critical for ensuring that outdated cache copies in the local cache are not utilized, thereby guaranteeing the coherence of DistR. Moreover, we repurpose one bit in the extension field to indicate whether the color has been updated. This design choice reduces the frequency of color updates during continuous local writes by the owner, thereby minimizing the frequency of color overflow.
@@ B.2 Immutable Reference
The procedure for dereferencing an immutable reference is outlined in Algorithm 4. This function accepts an immutable reference, consisting of a colored global address and a local copy address (see Figure 8), and outputs a local memory address for direct reading. The initial step involves determining the location of the object using the IsLocal function (Line 4). If this address is local, DistR clears the color embedded in the global address and returns it for direct access (Line 4).
For a remote object, DistR checks whether the local copy field is null (Line 4). If this field is null, indicating no recorded local cache copy, the algorithm proceeds to examine the cache hashmap $H$ on the server executing the access. This hashmap, keyed by the colored global address, contains entries pairing local copy addresses with a count of local immutable references pointing to this copy. The algorithm checks for the presence of the colored global address $r.g$ in $H$ (Line 4). If found, it increments the reference count in that entry and updates the local copy field in the immutable reference with the local address from the entry (Lines 4-4). If $r.g$ is not in $H$, the object is copied from the remote server, stored locally, and the local address is assigned to $r.l$ (Line 4). This newly fetched object, along with its associated reference count, is then added to $H$ (Line 4). Both the update and lookup operations on $H$ are performed atomically to avoid concurrency issues.
The procedure for dropping an immutable reference is outlined in the DropRef function (Lines 4-4) in Algorithm 4. It checks the local copy address field of the reference. If it is not null, DistR decrements the reference count for this local copy in $H$. Once the reference count hits zero, the local copy becomes eligible for eviction by DistR’s runtime system.
Figure 8: DistR’s pointer layout.
@@ 2 Function DerefMut($m$):
Algorithm 6 outlines the logic for dereferencing a mutable reference consisting of a colored global address and the owner’s address (see Figure 8). It returns a local memory address for writing. The first step in the algorithm is to determine the location of the referenced object (Line 6). If the object is in the local heap partition, DistR updates the color field in the mutable reference. This update is crucial for maintaining data integrity and coherence. A critical aspect of this mechanism is the first bit of the extension field, the U bit, which indicates whether the mutable reference’s color has been updated. The algorithm checks the U bit (Line 6); if it has not been updated, DistR increments the color field and sets the U bit to 1 (Lines 6-6) . This process prevents redundant color updates during subsequent dereferences, thereby enhancing efficiency.
For remote objects, DistR moves them to the local heap partition (Line 6). The move operation involves several steps: copying the object to a local address, updating the global address field in the mutable reference to this new address, and requesting the remote server to deallocate the original object.
Upon expiration of a mutable reference, as demonstrated in the DropMutRef function (Lines 6-6), DistR writes the mutable reference’s colored global address back to the owner’s global address field. This ensures that the owner maintains an up-to-date view of the data object. Additionally, writing back the updated color guarantees that data modifications through a mutable reference on the same machine as the owner can also be reflected in owner’s global address field. This reflection changes the key used for cache look-ups by subsequent immutable references, thereby preventing access to stale data copies cached on other servers.
@@ 1 Function DerefMut($p$, $\mathit{H}$):
In DistR, memory access through an owner can be mutable or immutable. The access type can be determined at runtime, but the presence of concurrent immutable references during immutable access is not always known. Therefore, similar to immutable references, the Box pointer for an owner employs a dual-address layout. Here we discuss the handling of both immutable (read) and mutable (write) access through the owner.
  • Immutable Access. As shown in Algorithm 7, this process begins with a check of the object location. If it is in local heap partition, the global address is directly returned for dereferencing (Lines 7-7). If not, DistR emulates the handling for immutable references (Lines 7-7). The algorithm checks if the local copy address field is null. If it is, DistR checks the hashmap for existing local copies, potentially copying data to local memory, updating the hashmap, and recording the local address in the local copy field. Otherwise, DistR directly dereferences this local address.
  • Mutable Access. As detailed in Algorithm 8, writing objects through the owner begins by checking the object’s location. If it is in local heap partition and the U bit is 0 (Line 8), the owner’s color will be updated and the U bit will be set (Lines 88). If the object is remote, instead of directly moving the object to local memory, this algorithm first checks the hashmap $H$ for a local copy (Line 8). If no local copy exists, DistR mimics the mutable reference dereferencing, moving the object to a local address and updating the owner’s global address field (Line 8). If a local copy is found, DistR retrieves this local address from $H$ and sets the global address field of the mutable reference to this local address, while also deallocating the original remote copy (Lines 88).
Furthermore, when an owner goes out of scope, DistR mimics the behavior of Rust’s standard dropping function for a Box pointer. It first drops every object owned by the referenced object, and then deallocates the object pointed to by this Box pointer. Importantly, the deallocation process in DistR during moving or dropping an object also asynchronously invalidates all cached copies of the object across all servers, ensuring that references to newly allocated objects at the same address do not retrieve stale cached data.
@@ Appendix C Memory Coherence Proof
We prove the memory coherence of DistR by demonstrating its adherence to two critical invariants as outlined in [57].
@@ C.1 Atomicity of Global-Address Change
In DistR, the colored global address is subject to modification in three specific scenarios:
@@ C.2 Global-Address-Change-on-Write Invariant
The complexity in this invariant arises from the inclusion of the U bit, which is employed as an optimization to reduce the frequency of color updates. As detailed at the end of § B.4, every time an immutable reference is created through the owner or a mutable reference, the U bit of that entity is reset to 0. Subsequently, when mutable dereferencing occurs on a mutable reference or Box pointer, the color in that entity will be changed and the U bit will be set to 1. Future mutable dereferencing through the same owner or mutable reference will not change color again until an immutable reference is borrowed from it. This optimization does not violate this Global-Address-Change-on-Write invariant, because access to data in Rust’s ownership model can be conceptually divided into epochs. Within each epoch, if immutable references exist, shared read access is permitted through multiple immutable references and the owner. Otherwise, the epoch allows only a single entity—either the owner or a mutable reference, but not both—to read or write data. Thus, modifications to the same data within the same epoch can be regarded as a singular modification. DistR ensures that at the end of each epoch, if the data has been altered, the colored global address will have changed. As for data accesses in a single epoch, if any entity modifies the data, subsequent reads can only occur through that modifier, which can directly read the updated data and does not involve the global address at all. Consequently, the Global-Address-Change-on-Write invariant still holds with our optimization after considering modifications within the same epoch as a single write operation.
@@ C.4 Stale-Value-Elimination Invariant
This invariant can be proved using the Global-Address-Change-on-Write invariant and the Updated-Global-Address-Visible invariant. As detailed in §C.2, data access can be categorized into different epochs. During each epoch, a value is either accessible through multiple immutable references or a single mutable reference. The value may be cached on various servers in epochs of shared immutable access. Consider a scenario where we are at the end of the N-th epoch, and the value undergoes modification in the subsequent (N+1)-th epoch. The Global-Address-Change-on-Write invariant ensures that the colored global address of this value is updated. Consequently, when the local cache hashmap is next queried in future epochs via an immutable reference to this value, it will reflect the updated global address, as dictated by the Updated-Global-Address-Visible invariant. This new colored global address will differ from any prior global addresses associated with this value, which serve as keys in each server’s local cache hashmaps. Therefore, no matching entries will be found in the hashmap, effectively preventing the reading of stale values in DistR.
Finally, the validity of the Data-Value invariant is established through a proof by contradiction. If the Data-Value invariant were violated, it would imply that a stale value is read post-mutation. The stale values only exist in the local read-only cache of each server, and values can only be mutated by the owner or mutable references. This indicates that a stale value copy stored in a server’s read-only cache is accessed again after the value is mutated through the owner or mutable references, which contradicts the Stale-Value-Elimination invariant. Hence, the Data-Value invariant holds true for DistR.
By establishing both the SWMR and Data-Value invariants for DistR, we have proved the correctness of our algorithm as well as the memory coherence of DistR.
@@ D.1 Borrowing Stack Values and Partial Borrowing
Rust allows for the borrowing of stack values and partial borrowing of struct components. This leads to scenarios where mutable references may exist to stack values or a part of a struct. The challenge in these cases is that their global addresses cannot be altered when accessed via a mutable reference on a different server. To resolve this, DistR employs a copy and write-back method. It involves copying the stack value or a part of the heap struct to local memory. Upon the expiration of the mutable reference, the modified local copy is written back to the original memory location on the original server. Additionally, DistR atomically increments the color field of the struct owner from which a partial borrow occurs. For stack values, lacking a Box pointer owner, an eager eviction policy is adopted for immutable references. When the reference count of a stack value drops to 0, we eagerly delete the local cache copy and its corresponding entry on that server, ensuring that subsequent references to the same stack value will read from its original location.
@@ D.2 Reference Creation and Ownership Transferring
During the creation of new shared immutable references from existing ones, only the global address field is duplicated, leaving the new reference’s local copy field null. This method maintains the integrity of the reference count in the hashmap and ensures that any local copy address is genuinely local. Following accesses through this new reference trigger updates to the local copy field and the hashmap by DistR.
In scenarios involving the transfer of ownership between servers, DistR deems it crucial to deallocate any local cache copies present on the source server. This step is essential to preserve consistency and prevent memory leaks during ownership transitions in distributed settings.

### v2

@@ A.3.2 Deploying DRust
Following configuration, build DRust as follows: {mdframed}[style=MyFrame,nobreak=true] # Compile the communication static library
@@
# Copy the static library to the DRust directory
@@
cargo build ––release
@@ A.3.3 Running Applications
DRust is bundled with four example applications: Dataframe, GEMM, KVStore, and SocialNet. Follow these steps to execute them:
@@
    Launch the DRust executable on all servers, excluding the main server: {mdframed}[style=MyFrame,nobreak=true] # Start the DRust process with the specified server index and application name.
@@
    On the main server: {mdframed}[style=MyFrame,nobreak=true] # Start the main DRust process with the specified application.
@@
More details of DRust’s installation and deployment can be found in DRust’s code repository.
@@ B.1 Pointer Layout
DRust extends each Rust Box pointer and reference with an additional 64-bit field, as depicted in Figure 8. Additionally, in the global address field, we reserve 16 bits as “color” bits, which are used to track the version of the referenced value. Every time a mutable reference is dereferenced for the first time, or the owner performs a write operation subsequent to the return of all immutable references, the color bits are updated. Consequently, this mechanism guarantees that any newly created immutable references will obtain a new global address [^3] In the following text, a “colored global address” means the composition of a global address and its color.. This update is critical for ensuring that outdated cache copies in the local cache are not utilized, thereby guaranteeing the coherence of DRust. Moreover, we repurpose one bit in the extension field to indicate whether the color has been updated. This design choice reduces the frequency of color updates during continuous local writes by the owner, thereby minimizing the frequency of color overflow.
@@ B.2 Immutable Reference
The procedure for dereferencing an immutable reference is outlined in Algorithm 4. This function accepts an immutable reference, consisting of a colored global address and a local copy address (see Figure 8), and outputs a local memory address for direct reading. The initial step involves determining the location of the object using the IsLocal function (Line 4). If this address is local, DRust clears the color embedded in the global address and returns it for direct access (Line 4).
For a remote object, DRust checks whether the local copy field is null (Line 4). If this field is null, indicating no recorded local cache copy, the algorithm proceeds to examine the cache hashmap $H$ on the server executing the access. This hashmap, keyed by the colored global address, contains entries pairing local copy addresses with a count of local immutable references pointing to this copy. The algorithm checks for the presence of the colored global address $r.g$ in $H$ (Line 4). If found, it increments the reference count in that entry and updates the local copy field in the immutable reference with the local address from the entry (Lines 4-4). If $r.g$ is not in $H$, the object is copied from the remote server, stored locally, and the local address is assigned to $r.l$ (Line 4). This newly fetched object, along with its associated reference count, is then added to $H$ (Line 4). Both the update and lookup operations on $H$ are performed atomically to avoid concurrency issues.
The procedure for dropping an immutable reference is outlined in the DropRef function (Lines 4-4) in Algorithm 4. It checks the local copy address field of the reference. If it is not null, DRust decrements the reference count for this local copy in $H$. Once the reference count hits zero, the local copy becomes eligible for eviction by DRust’s runtime system.
Figure 8: DRust’s pointer layout.
@@ 2 Function DerefMut($m$):
Algorithm 6 outlines the logic for dereferencing a mutable reference consisting of a colored global address and the owner’s address (see Figure 8). It returns a local memory address for writing. The first step in the algorithm is to determine the location of the referenced object (Line 6). If the object is in the local heap partition, DRust updates the color field in the mutable reference. This update is crucial for maintaining data integrity and coherence. A critical aspect of this mechanism is the first bit of the extension field, the U bit, which indicates whether the mutable reference’s color has been updated. The algorithm checks the U bit (Line 6); if it has not been updated, DRust increments the color field and sets the U bit to 1 (Lines 6-6) . This process prevents redundant color updates during subsequent dereferences, thereby enhancing efficiency.
For remote objects, DRust moves them to the local heap partition (Line 6). The move operation involves several steps: copying the object to a local address, updating the global address field in the mutable reference to this new address, and requesting the remote server to deallocate the original object.
Upon expiration of a mutable reference, as demonstrated in the DropMutRef function (Lines 6-6), DRust writes the mutable reference’s colored global address back to the owner’s global address field. This ensures that the owner maintains an up-to-date view of the data object. Additionally, writing back the updated color guarantees that data modifications through a mutable reference on the same machine as the owner can also be reflected in owner’s global address field. This reflection changes the key used for cache look-ups by subsequent immutable references, thereby preventing access to stale data copies cached on other servers.
@@ 1 Function DerefMut($p$, $\mathit{H}$):
In DRust, memory access through an owner can be mutable or immutable. The access type can be determined at runtime, but the presence of concurrent immutable references during immutable access is not always known. Therefore, similar to immutable references, the Box pointer for an owner employs a dual-address layout. Here we discuss the handling of both immutable (read) and mutable (write) access through the owner.
  • Immutable Access. As shown in Algorithm 7, this process begins with a check of the object location. If it is in local heap partition, the global address is directly returned for dereferencing (Lines 7-7). If not, DRust emulates the handling for immutable references (Lines 7-7). The algorithm checks if the local copy address field is null. If it is, DRust checks the hashmap for existing local copies, potentially copying data to local memory, updating the hashmap, and recording the local address in the local copy field. Otherwise, DRust directly dereferences this local address.
  • Mutable Access. As detailed in Algorithm 8, writing objects through the owner begins by checking the object’s location. If it is in local heap partition and the U bit is 0 (Line 8), the owner’s color will be updated and the U bit will be set (Lines 88). If the object is remote, instead of directly moving the object to local memory, this algorithm first checks the hashmap $H$ for a local copy (Line 8). If no local copy exists, DRust mimics the mutable reference dereferencing, moving the object to a local address and updating the owner’s global address field (Line 8). If a local copy is found, DRust retrieves this local address from $H$ and sets the global address field of the mutable reference to this local address, while also deallocating the original remote copy (Lines 88).
Furthermore, when an owner goes out of scope, DRust mimics the behavior of Rust’s standard dropping function for a Box pointer. It first drops every object owned by the referenced object, and then deallocates the object pointed to by this Box pointer. Importantly, the deallocation process in DRust during moving or dropping an object also asynchronously invalidates all cached copies of the object across all servers, ensuring that references to newly allocated objects at the same address do not retrieve stale cached data.
@@ Appendix C Memory Coherence Proof
We prove the memory coherence of DRust by demonstrating its adherence to two critical invariants as outlined in [57].
@@ C.1 Atomicity of Global-Address Change
In DRust, the colored global address is subject to modification in three specific scenarios:
@@ C.2 Global-Address-Change-on-Write Invariant
The complexity in this invariant arises from the inclusion of the U bit, which is employed as an optimization to reduce the frequency of color updates. As detailed at the end of § B.4, every time an immutable reference is created through the owner or a mutable reference, the U bit of that entity is reset to 0. Subsequently, when mutable dereferencing occurs on a mutable reference or Box pointer, the color in that entity will be changed and the U bit will be set to 1. Future mutable dereferencing through the same owner or mutable reference will not change color again until an immutable reference is borrowed from it. This optimization does not violate this Global-Address-Change-on-Write invariant, because access to data in Rust’s ownership model can be conceptually divided into epochs. Within each epoch, if immutable references exist, shared read access is permitted through multiple immutable references and the owner. Otherwise, the epoch allows only a single entity—either the owner or a mutable reference, but not both—to read or write data. Thus, modifications to the same data within the same epoch can be regarded as a singular modification. DRust ensures that at the end of each epoch, if the data has been altered, the colored global address will have changed. As for data accesses in a single epoch, if any entity modifies the data, subsequent reads can only occur through that modifier, which can directly read the updated data and does not involve the global address at all. Consequently, the Global-Address-Change-on-Write invariant still holds with our optimization after considering modifications within the same epoch as a single write operation.
@@ C.4 Stale-Value-Elimination Invariant
This invariant can be proved using the Global-Address-Change-on-Write invariant and the Updated-Global-Address-Visible invariant. As detailed in §C.2, data access can be categorized into different epochs. During each epoch, a value is either accessible through multiple immutable references or a single mutable reference. The value may be cached on various servers in epochs of shared immutable access. Consider a scenario where we are at the end of the N-th epoch, and the value undergoes modification in the subsequent (N+1)-th epoch. The Global-Address-Change-on-Write invariant ensures that the colored global address of this value is updated. Consequently, when the local cache hashmap is next queried in future epochs via an immutable reference to this value, it will reflect the updated global address, as dictated by the Updated-Global-Address-Visible invariant. This new colored global address will differ from any prior global addresses associated with this value, which serve as keys in each server’s local cache hashmaps. Therefore, no matching entries will be found in the hashmap, effectively preventing the reading of stale values in DRust.
Finally, the validity of the Data-Value invariant is established through a proof by contradiction. If the Data-Value invariant were violated, it would imply that a stale value is read post-mutation. The stale values only exist in the local read-only cache of each server, and values can only be mutated by the owner or mutable references. This indicates that a stale value copy stored in a server’s read-only cache is accessed again after the value is mutated through the owner or mutable references, which contradicts the Stale-Value-Elimination invariant. Hence, the Data-Value invariant holds true for DRust.
By establishing both the SWMR and Data-Value invariants for DRust, we have proved the correctness of our algorithm as well as the memory coherence of DRust.
@@ D.1 Borrowing Stack Values and Partial Borrowing
Rust allows for the borrowing of stack values and partial borrowing of struct components. This leads to scenarios where mutable references may exist to stack values or a part of a struct. The challenge in these cases is that their global addresses cannot be altered when accessed via a mutable reference on a different server. To resolve this, DRust employs a copy and write-back method. It involves copying the stack value or a part of the heap struct to local memory. Upon the expiration of the mutable reference, the modified local copy is written back to the original memory location on the original server. Additionally, DRust atomically increments the color field of the struct owner from which a partial borrow occurs. For stack values, lacking a Box pointer owner, an eager eviction policy is adopted for immutable references. When the reference count of a stack value drops to 0, we eagerly delete the local cache copy and its corresponding entry on that server, ensuring that subsequent references to the same stack value will read from its original location.
@@ D.2 Reference Creation and Ownership Transferring
During the creation of new shared immutable references from existing ones, only the global address field is duplicated, leaving the new reference’s local copy field null. This method maintains the integrity of the reference count in the hashmap and ensures that any local copy address is genuinely local. Following accesses through this new reference trigger updates to the local copy field and the hashmap by DRust.
In scenarios involving the transfer of ownership between servers, DRust deems it crucial to deallocate any local cache copies present on the source server. This step is essential to preserve consistency and prevent memory leaks during ownership transitions in distributed settings.